# Ultra-cold atoms as quantum simulators for relativistic phenomena

### v1

@@ 2.2.2   Hawking radiation as tunneling?
To conclude this section, let us discuss another frequently used interpretation of Hawking radiation as a tunneling process, see, e.g., [125]. Clearly, there are similarities: Both are quantum effects which generate a tail at finite spatial distances (the evanescent part of the wave function in tunneling and the correlation of the vacuum fluctuations of quantum fields) which does not exist classically. However, when picturing Hawking radiation as tunneling of a particle from the interior of the black hole to the exterior, one should also be aware of some crucial differences.
@@
Let us start by considering the case of relativistic quantum fields without modifications of the dispersion relation – which is the case discussed in most of the papers devoted to the tunneling picture. First, tunneling in quantum mechanics usually refers the a case where, in classical mechanics, there is no trajectory going from a point in front of the potential barrier to a point behind the barrier if the initial kinetic energy is smaller than the barrier height. Yet, in quantum mechanics, the propagator is non-vanishing (though exponentially suppressed). In contrast, in relativistic quantum field theory, the causal propagator from any space-time point inside the black hole (i.e., behind the horizon) to any space-time point outside vanishes exactly due to causality (without any exponential tails which would correspond to tunneling in quantum mechanics). Second, as a related point, the WKB phase displays a branch-cut singularity at the classical turning points for tunneling – but this singularity is resolved when calculating the full wave function (i.e., beyond the WKB approximation) in quantum mechanics. The situation is different for Hawking radiation in relativistic quantum field theory since there the stationary mode functions at the horizon acquire a different kind of phase singularity which does not go away when considering the full wave equation. Actually, this phase singularity is closely related to the trans-Planckian problem mentioned above. Since the tunneling picture of Hawking radiation makes heavy use of these singularities, see, e.g., [125], this difference is also quite important [^10] Probably one reason for the relative popularity of the tunneling picture for Hawking radiation is that it allows a “short-cut derivation” of the Hawking temperature without a full-fledged quantum-field theory calculation. However, in view of the complexity of the problem, it is important to remember that such a short cut can slur over some of the subtleties of the derivation – which is the main motivation for this subsection.. Third, for tunneling in quantum mechanics, the state which comes out at the rear end of the barrier depends on the state (energy and amplitude etc.) which was incident on the front end (i.e., ”trying” to tunnel through). The outgoing Hawking radiation, on the other hand, is independent of the quantum state inside the black hole (at least within relativistic quantum field theory). If there was such a simple connection between the quantum state inside the black hole and the outgoing Hawking radiation as in ordinary tunneling in quantum mechanics, it would have profound consequences for our discussions of the black-hole information puzzle. Fourth, for the Sauter-Schwinger effect discussed in Sec. 2.8, the Dirac equation in the presence of a constant electric field, for example, can be cast into a tunneling problem with a spatially oscillating wave function outside the potential barrier and an exponentially decaying wave function in between the classical turning points (i.e., inside the barrier). Thus, in the case of the Sauter-Schwinger effect, the interpretation as tunneling (from the Dirac sea to the positive continuum) can be justified qualitatively as well as quantitatively in this way. For Hawking radiation, on the other hand, I am not aware of such a mapping and it is also not clear what the two classical turning points should be in this case. In view of the points above, such a direct mapping probably does not exist.
@@
Note that, already in the original paper [125], it has been mentioned that an alternative interpretation could be to replace the tunneling of the Hawking particle from inside to outside by the tunneling of its partner particle into the black hole – thereby liberating the Hawking particle to go off to infinity as a real particle. This alternative interpretation evades some of the problems mentioned above and also incorporates the important point of connecting the outgoing Hawking particles with their in-falling partners – which was actually important for the Technion experiment mentioned above.
@@ 2.3   Unruh effect
Via the principle of equivalence, the Unruh effect is closely related to Hawking radiation. Actually, Unruh discovered it when trying to understand Hawking radiation better [128]. However, in order to explain this effect, let us first go to flat space-time and consider the usual Minkowski vacuum. Then, imagine placing a photon detector at rest in this vacuum state. Obviously (neglecting possible dark counts) the detector does not click. If we now imagine moving this detector with a constant velocity [^12] Which is of course smaller than the speed of light, see Sec. 2.6. then it still does not click because the Minkowski vacuum is Lorentz invariant. However, the Unruh effect predicts that the situation changes if we accelerate the detector. In the case of uniform acceleration $a$, the detector responds as if it was immersed in a thermal bath with the Unruh temperature
@@
Now let us come back to the relation to Hawking radiation via the principle of equivalence as mentioned above. To this end, let us consider a series of static observers at fixed distances to the black-hole horizon and compare their findings with that of a freely falling observer which starts with small velocity far away from the black hole and then falls in. At large distances, the static and the freely falling observers both see basically the same Hawking radiation emitted by the black hole. However, this agreement changes when the freely falling observer approaches the black hole. On the one hand – as we have already discussed above – the freely falling observer does not notice anything special at the horizon but locally sees a vacuum state. On the other hand, in view of the gravitational red-shift of the photons propagating away from the black hole, static observers closer to the horizon must find a higher temperature than static observers further away for their observations to be consistent (in a stationary quantum state). Thus, the closer they are to the horizon (from which the gravitational red-shift becomes infinitely strong) the higher their observed temperatures must be. This explains the connection to the Unruh effect since these static observers feel a strong gravitational pull and thus must be strongly accelerated (in view of the principle of equivalence) in order to retain their fixed distance to the horizon. In contrast, the freely falling observer corresponds locally to the inertial Minkowski observer. The latter does not detect any photons while the former sees a thermal bath.
@@ 2.3.1   Analogue of Unruh effect
Now, another important ingredient for realizing an analogue of the Unruh effect is the detector. Instead of photons, this detector should detect phonons, i.e., density and phase fluctuations of the Bose-Einstein condensate. One promising option could be an atomic quantum dot, i.e., a very tight optical potential (like optical tweezers) which may hold one or a small number of atoms, see, e.g., [132, 30]. The internal states of this detector could then correspond to the internal states of the trapped atoms or their motional states or even to the number of trapped atoms inside the atomic quantum dot. The atomic quantum dot could be coupled to remaining Bose-Einstein condensate via tunneling or optical transitions, possibly in connection with Rydberg couplings or dipolar couplings.
@@
In order to simulate the accelerated detector, one could imagine moving the atomic quantum dot through the condensate on a non-inertial trajectory, see, e.g., [133]. This would model the coupling of the detector to the field along such a non-inertial trajectory. However, another important ingredient for the Unruh effect is the relativistic time dilatation of the detector (i.e, the difference between the Minkowski time $t$ and the proper time $\tau$ of the detector), which is not represented by the atomic quantum dot a priori. To take this effect into account, one could externally tune the internal eigen-energies of the atomic quantum dot such that they represent the proper time $\tau$. Alternatively, one could replace the uniformly accelerated trajectory by a circular motion of the detector, e.g., $x(t)=R\cos(\Omega t)$, $y(t)=R\sin(\Omega t)$ and $z=0$. In this case, the absolute value of the acceleration would be constant but its direction would rotate constantly. The two-point function along the detector trajectory can be obtained by inserting $x(t)=R\cos(\Omega t)$, $y(t)=R\sin(\Omega t)$ and $z=0$ into Eq. (2.16). The result would be stationary, i.e., only depend on $t-t^{\prime}$, where this dependence is periodic in real time – but no longer in imaginary time. As a consequence, while the detector on a circular trajectory would also experience particles, their spectrum would not be exactly thermal anymore. On the other hand, one would not have to tune the internal eigen-energies of the atomic quantum since the Lorentz factor determining the relativistic time dilatation of the detector is constant for the circular motion. In principle, this non-thermal excitation spectrum could be measured by moving the atomic quantum dot through the condensate on a circular trajectory, see, e.g., [133].
@@
However, in order to detect the analogue of the Unruh effect as a quantum phenomenon, one should be careful to avoid creating classical excitations by stirring the condensate with the moving atomic quantum dot – which then could also excite it. This is often a problem when trying to observe quantum effects, as competing classical background effects can easily be larger than the usually small quantum phenomena one is interested in. As an idea to avoid these problems, it has been suggested [134] to replace the atomic quantum dots by laser beams which measure the local density of the condensate and move through it on circular orbits, see also [135]. By analysing the density fluctuations measured by those laser beams (in an interferometric set-up), one could then reconstruct the detector response. Using two laser beams, one red and one blue detuned from the resonance frequency of the condensate atoms, one can suppress the classical back-reaction of those laser beams onto the condensate, i.e., avoid unwanted excitations. However, one should still be careful because every measurement perturbs the quantum state of the system – which might in turn react back on the density fluctuations. This back-reaction effect could be minimized by using weak instead of strong measurements of the density fluctuations.
@@
There have also been proposals which focus on other aspects of the Unruh effect instead of the non-inertial trajectories. For example, in [136, 137] an optical lattice simulator for the Rindler geometry (i.e., the Minkowski metric in terms of the Rindler coordinates) has been proposed. In order to observe particle creation, the idea is to start in the ground state of the Minkowski metric and then to suddenly switch (i.e., quench) to the Rindler geometry. This scenario displays some similarities to the (sudden) collapse to a black hole, as depicted in Fig. 2.1, and also to the phenomenon of cosmological particle creation as discussed in the next section. Other proposals focus on the squeezing mechanism because the Minkoswki vacuum is a squeezed state in terms of the Rindler creation and annihilation operators (associated to the uniformly accelerated observers). However, since squeezing is important for many phenomena (including Hawking radiation as discussed in the previous section as well as cosmological particle creation as discussed in the next section), this feature alone is probably not sufficient for claiming to have a simulator for the Unruh effect. As in all of these simulators, it is important to carefully determine which aspects of the phenomenon to be simulated are reproduced and which not – or, equivalently, how many and which steps lie between the original phenomenon and the simulator.
@@ 2.4   Cosmological particle creation
Already in the early days of quantum mechanics (long before Hawking and Unruh etc.), Schrödinger [138] understood that the expansion of the Universe can create particles out of the vacuum, see also [139]. In principle, this phenomenon can also be traced back to the tearing apart process as already discussed in Sec. 2.2 but now this tearing apart is caused by the time-dependent expansion of the Universe instead of the stationary gravitational field of the black hole. However, one may also understand the basic mechanism by means of a single harmonic oscillator. To show this, let us consider the spatially flat Friedmann-Lemaitre-Robertson-Walker metric [^13] In case of spatial curvature, the spatial line element (which reads $d\mbox{\boldmath$r$}^{2}=dx^{2}+dy^{2}+dz^{2}$ in 3+1 dimensions) would acquire additional position dependent metric coefficients, as discussed below.
@@
$\left(\frac{d^{2}}{dT^{2}}-a^{2(D-1)}\mbox{\boldmath$k$}^{2}\right)\phi_{\mbox{\boldmath$\scriptstyle k$}}=0\,,$ (2.19)
@@
Now, in analogy to the line of arguments in Sec. 2.2, if we start in the ground state of this harmonic oscillator and change its frequency $\Omega_{\mbox{\boldmath$\scriptstyle k$}}(T)$ slowly enough (i.e., $\dot{\Omega}_{\mbox{\boldmath$\scriptstyle k$}}\ll\Omega^{2}_{\mbox{\boldmath$\scriptstyle k$}}$ and $\ddot{\Omega}_{\mbox{\boldmath$\scriptstyle k$}}\ll\Omega^{3}_{\mbox{\boldmath$\scriptstyle k$}}$ etc.) the quantum evolution will be adiabatic, i.e., the quantum state will stay close to the instantaneous ground state in view of the adiabatic theorem. However, if the time-dependence of $\Omega_{\mbox{\boldmath$\scriptstyle k$}}(T)$ is too fast, there will be deviations, i.e., the final state will differ from the ground state – it will be a squeezed state. As an extreme example, let us consider a sudden change of $\Omega_{\mbox{\boldmath$\scriptstyle k$}}(T)$ which is usually referred to as a quench. For such a sudden switch, the wave function has no time to change, i.e., immediately after the quench it will have the same form as before, see Fig. 2.4. For the initial value $\Omega_{\mbox{\boldmath$\scriptstyle k$}}^{\rm in}$ this wave function represents the ground state and thus does not change with time. After the quench to $\Omega_{\mbox{\boldmath$\scriptstyle k$}}^{\rm out}\neq\Omega_{\mbox{\boldmath$\scriptstyle k$}}^{\rm in}$, however, this wave function has the “wrong” width and thus does not represent the ground state for the new value of $\Omega_{\mbox{\boldmath$\scriptstyle k$}}^{\rm out}$. Instead, it is an excited state – in this case, a squeezed state containing pairs of excitation quanta. As a result, it is also no longer a stationary state, but it starts to “breathe” by changing its width in an oscillatory manner, see Fig. 2.4.

### v2

@@ 2.2.2   Hawking radiation as tunneling?
To conclude this section, let us discuss another frequently used interpretation of Hawking radiation as a tunneling process, see, e.g., [127]. Clearly, there are similarities: Both are quantum effects which generate a tail at finite spatial distances (the evanescent part of the wave function in tunneling and the correlation of the vacuum fluctuations of quantum fields) which does not exist classically. However, when picturing Hawking radiation as tunneling of a particle from the interior of the black hole to the exterior in complete analogy to tunneling through a potential barrier in ordinary quantum mechanics, one should also be aware of some crucial differences.
@@
Let us start by considering the case of relativistic quantum fields without modifications of the dispersion relation – which is the case discussed in most of the papers devoted to the tunneling picture. First, tunneling in quantum mechanics usually refers the a case where, in classical mechanics, there is no trajectory going from a point in front of the potential barrier to a point behind the barrier if the initial kinetic energy is smaller than the barrier height. Yet, in quantum mechanics, the propagator is non-vanishing (though exponentially suppressed). In contrast, in relativistic quantum field theory, the causal propagator from any space-time point inside the black hole (i.e., behind the horizon) to any space-time point outside vanishes exactly [^13] Note that the causal propagator is (already in flat space-time) different from the two-point (or Wightman) function. The latter can be non-zero for space-like separated points, which reflects the quantum vacuum correlations, but does not correspond to real propagation. due to causality (without any exponential tails which would correspond to tunneling in ordinary quantum mechanics). Second, as a related point, the WKB phase displays a branch-cut singularity at the classical turning points for tunneling – but this singularity is resolved when calculating the full wave function (i.e., beyond the WKB approximation) in quantum mechanics. The situation is different for Hawking radiation in relativistic quantum field theory since there the stationary mode functions at the horizon acquire a different kind of phase singularity which does not go away when considering the full wave equation. Actually, this phase singularity is closely related to the trans-Planckian problem mentioned above. Since the tunneling picture of Hawking radiation makes heavy use of these singularities, see, e.g., [127], this difference is also quite important [^14] Probably one reason for the relative popularity of the tunneling picture for Hawking radiation is that it allows a “short-cut derivation” of the Hawking temperature without a full-fledged quantum-field theory calculation. However, in view of the complexity of the problem, it is important to remember that such a short cut can gloss over some of the subtleties of the derivation – which is one of the the main motivations for this subsection.. Third, for tunneling in ordinary quantum mechanics, the state which comes out at the rear end of the barrier depends on the state (energy and amplitude etc.) which was incident on the front end (i.e., ”trying” to tunnel through). The outgoing Hawking radiation, on the other hand, is independent of the quantum state inside the black hole (at least within relativistic quantum field theory). If there was such a simple connection between the quantum state inside the black hole and the outgoing Hawking radiation as in ordinary tunneling in quantum mechanics, it would have profound consequences for our discussions of the black-hole information puzzle. Fourth, for the Sauter-Schwinger effect discussed in Sec. 2.8, the Dirac equation in the presence of a constant electric field, for example, can be cast into a tunneling problem with a spatially oscillating wave function outside the potential barrier and an exponentially decaying wave function in between the classical turning points (i.e., inside the barrier). Thus, in the case of the Sauter-Schwinger effect, the interpretation as tunneling (from the Dirac sea to the positive continuum) can be justified qualitatively as well as quantitatively in this way. For Hawking radiation, on the other hand, I am not aware of such a mapping and it is also not clear what the two classical turning points should be in this case. In view of the points above, such a direct mapping probably does not exist.
As a disclaimer, it should be stressed here that the above arguments apply to tunneling in real space, i.e., from the interior to the exterior of the black hole. For tunneling in a different space – such as in configuration space – these arguments should be revisited. For example, it could well be that, after we understand Hawking radiation based on a full theory of quantum gravity, we shall find that it can be interpreted as tunneling from one configuration (of the coupled system space-time plus matter) to another one. In this case, the above arguments suggest that this tunneling process occurring in an abstract configuration space cannot be simply projected onto tunneling of matter from the interior to the exterior of the black hole, i.e., in real space.
@@
Also note that, already in the original paper by Parikh and Wilczek [127], it has been mentioned that an alternative interpretation could be to replace the tunneling of the Hawking particle from inside to outside by the tunneling of its partner particle into the black hole – thereby liberating the Hawking particle to go off to infinity as a real particle. This alternative interpretation has already been described by Hawking in his seminal paper [109] and evades some of the problems mentioned above. It also incorporates the important point of connecting the outgoing Hawking particles with their in-falling partners – which was actually important for the Technion experiment mentioned above.
@@ 2.3   Unruh effect
Via the principle of equivalence, the Unruh effect is closely related to Hawking radiation. Actually, Unruh discovered it [130] basically simultaneously with the prediction of black-hole evaporation [108, 109] and then realized the relation between the two phenomena. However, in order to explain the Unruh effect, let us first go to flat space-time and consider the usual Minkowski vacuum. Then, imagine placing a photon detector at rest in this vacuum state. Obviously (neglecting possible dark counts) the detector does not click. If we now imagine moving this detector with a constant velocity [^16] Which is of course smaller than the speed of light, see Sec. 2.6. then it still does not click because the Minkowski vacuum is Lorentz invariant. However, the Unruh effect predicts that the situation changes if we accelerate the detector. In the case of uniform acceleration $a$, the detector responds as if it was immersed in a thermal bath with the Unruh temperature
@@
Now let us come back to the relation to Hawking radiation via the principle of equivalence as mentioned above. To this end, let us consider a series of static observers at fixed distances to the black-hole horizon and compare their findings with that of a freely falling observer which starts with small velocity far away from the black hole and then falls in. At large distances, the static and the freely falling observers do not differ much and thus both see basically the same Hawking radiation emitted by the black hole. However, this agreement changes when the freely falling observer approaches the black hole. On the one hand – as we have already discussed above – the freely falling observer does not notice anything special at the horizon but locally sees a vacuum state. On the other hand, in view of the gravitational red-shift of the photons propagating away from the black hole, static observers closer to the horizon must find a higher temperature than static observers further away for their observations to be consistent (in a stationary quantum state). Thus, the closer the static observers are to the horizon (from which the gravitational red-shift becomes infinitely strong) the higher their observed temperatures must be. This explains the connection to the Unruh effect since these static observers feel a strong gravitational pull and thus must be strongly accelerated (in view of the principle of equivalence) in order to retain their fixed distance to the horizon. In contrast, the freely falling observer corresponds locally to the inertial Minkowski observer. The latter does not detect any photons while the former sees a thermal bath.
@@ 2.3.1   Analogue of Unruh effect
Now, another important ingredient for realizing an analogue of the Unruh effect is the detector. Instead of photons, this detector should detect phonons, i.e., density and phase fluctuations of the Bose-Einstein condensate. One promising option could be an atomic quantum dot, i.e., a very tight optical potential (like optical tweezers) which may hold one or a small number of atoms, see, e.g., [134, 135, 136, 30]. The internal states of this detector could then correspond to the internal states of the trapped atoms or their motional states or even to the number of trapped atoms inside the atomic quantum dot. The atomic quantum dot could be coupled to remaining Bose-Einstein condensate via tunneling or optical transitions, possibly in connection with Rydberg couplings or dipolar couplings.
@@
In order to simulate the accelerated detector, one could imagine moving the atomic quantum dot through the condensate on a non-inertial trajectory, see, e.g., [137]. This would model the coupling of the detector to the field along such a non-inertial trajectory. However, another important ingredient for the Unruh effect is the relativistic time dilatation of the detector (i.e, the difference between the Minkowski time $t$ and the proper time $\tau$ of the detector), which is not represented by the atomic quantum dot a priori. To take this effect into account, one could externally tune the internal eigen-energies of the atomic quantum dot such that they represent the proper time $\tau$. Alternatively, one could replace the uniformly accelerated trajectory by a circular motion of the detector, e.g., $x(t)=R\cos(\Omega t)$, $y(t)=R\sin(\Omega t)$ and $z=0$. In this case, the absolute value of the acceleration would be constant but its direction would rotate constantly. The two-point function along the detector trajectory can be obtained by inserting $x(t)=R\cos(\Omega t)$, $y(t)=R\sin(\Omega t)$ and $z=0$ into Eq. (2.16). The result would be stationary, i.e., only depend on $t-t^{\prime}$, but this dependence is no longer periodic in imaginary time. As a consequence, while the detector on a circular trajectory would also experience particles, their spectrum would not be exactly thermal anymore. On the other hand, one would not have to tune the internal eigen-energies of the atomic quantum since the Lorentz factor determining the relativistic time dilatation of the detector is constant for the circular motion. In principle, this non-thermal excitation spectrum could be measured by moving the atomic quantum dot through the condensate on a circular trajectory, see, e.g., [137].
@@
However, in order to detect the analogue of the Unruh effect as a quantum phenomenon, one should be careful to avoid creating classical excitations by stirring the condensate with the moving atomic quantum dot – which then could also excite it. This is often a problem when trying to observe quantum effects, as competing classical background effects can easily be larger than the usually small quantum phenomena one is interested in. As an idea to avoid these problems, it has been suggested [138] to replace the atomic quantum dots by laser beams which measure the local density of the condensate and move through it on circular orbits, see also [139, 140]. By analysing the density fluctuations measured by those laser beams (in an interferometric set-up), one could then reconstruct the detector response. Using two laser beams, one red and one blue detuned from the resonance frequency of the condensate atoms, one can suppress the classical back-reaction of those laser beams onto the condensate, i.e., avoid unwanted excitations. However, one should still be careful because every measurement perturbs the quantum state of the system – which might in turn react back on the density fluctuations. This back-reaction effect could be minimized by using weak instead of strong measurements of the density fluctuations.
@@
There have also been proposals which focus on other aspects of the Unruh effect instead of the non-inertial trajectories. For example, in [141, 142] an optical lattice simulator for the Rindler geometry (i.e., the Minkowski metric in terms of the Rindler coordinates) has been proposed. In order to observe particle creation, the idea is to start in the ground state of the Minkowski metric and then to suddenly switch (i.e., quench) to the Rindler geometry, see also [143]. This scenario displays some similarities to the (sudden) collapse to a black hole, as depicted in Fig. 2.1, and also to the phenomenon of cosmological particle creation as discussed in the next section. Other proposals focus on the squeezing mechanism because the Minkoswki vacuum is a squeezed state in terms of the Rindler creation and annihilation operators (associated to the uniformly accelerated observers). However, since squeezing is important for many phenomena (including Hawking radiation as discussed in the previous section as well as cosmological particle creation as discussed in the next section), this feature alone is probably not sufficient for claiming to have a simulator for the Unruh effect. As in all of these simulators, it is important to carefully determine which aspects of the phenomenon to be simulated are reproduced and which not – or, equivalently, how many and which steps lie between the original phenomenon and the simulator.
@@ 2.4   Cosmological particle creation
Already in the early days of quantum mechanics (long before Hawking and Unruh etc.), Schrödinger [144] understood that the expansion of the Universe can create particles out of the vacuum, see also [145]. In principle, this phenomenon can also be traced back to the tearing apart process as already discussed in Sec. 2.2 but now this tearing apart is caused by the time-dependent expansion of the Universe instead of the stationary gravitational field of the black hole (long after the gravitational collapse). However, one may also understand the basic mechanism by means of a single harmonic oscillator. To show this, let us consider the spatially flat Friedmann-Lemaitre-Robertson-Walker metric [^17] In case of spatial curvature, the spatial line element (which reads $d\mbox{\boldmath$r$}^{2}=dx^{2}+dy^{2}+dz^{2}$ in 3+1 dimensions) would acquire additional position dependent metric coefficients, as discussed below.
@@
${\left(\frac{d^{2}}{dT^{2}}+a^{2(D-1)}\mbox{\boldmath$k$}^{2}\right)\phi_{\mbox{\boldmath$\scriptstyle k$}}=0}\,,$ (2.19)
@@
Now, in analogy to the line of arguments in Sec. 2.2, if we start in the ground state of this harmonic oscillator and change its frequency $\Omega_{\mbox{\boldmath$\scriptstyle k$}}(T)$ slowly enough (i.e., $\dot{\Omega}_{\mbox{\boldmath$\scriptstyle k$}}\ll\Omega^{2}_{\mbox{\boldmath$\scriptstyle k$}}$ and $\ddot{\Omega}_{\mbox{\boldmath$\scriptstyle k$}}\ll\Omega^{3}_{\mbox{\boldmath$\scriptstyle k$}}$ etc.) the quantum evolution will be adiabatic, i.e., the quantum state will stay close to the instantaneous ground state in view of the adiabatic theorem. However, if the time-dependence of $\Omega_{\mbox{\boldmath$\scriptstyle k$}}(T)$ is too fast, there will be deviations, i.e., the final state will differ from the ground state – it will be a squeezed state. As an extreme example, let us consider a sudden change of $\Omega_{\mbox{\boldmath$\scriptstyle k$}}(T)$ which is usually referred to as a quench. For such a sudden switch, the wave function has no time to change, i.e., immediately after the quench it will have the same form as before, see Fig. 2.4. For the initial value $\Omega_{\mbox{\boldmath$\scriptstyle k$}}^{\rm in}$ this wave function represents the ground state and thus does not change with time. After the quench to $\Omega_{\mbox{\boldmath$\scriptstyle k$}}^{\rm out}\neq\Omega_{\mbox{\boldmath$\scriptstyle k$}}^{\rm in}$, however, this wave function has the “wrong” width and thus does not represent the ground state for the new value of $\Omega_{\mbox{\boldmath$\scriptstyle k$}}^{\rm out}$. Instead, it is an excited state – or, more precisely, a superposition state of the ground state and (even) excited states – in this case, a squeezed state containing pairs of excitation quanta. As a result, it is also no longer a stationary state, but it starts to “breathe” by changing its width in an oscillatory manner, see Fig. 2.4.